ELCTRON WITH INITIAL K.E 100ev is accelerated through a P.D. OF 50V NOW de Brogile WAVELENTH BECOMES WHAT

Grade:

2 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Pushpak,

kinetic energy after acceleration=100+50=150eV, let E be energy,p be momentum,c be speed of light in vacuum and Me be rest mass of electron. E=150eV

E2= p2c2+m2c4, from this equation find p,λ=h/p

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Pushpak !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR

Aman Bansal
12 years ago

Dear Pushpak,

Using the above formula and using enery relation we can evaluate de broglies wavelength.

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert